OpenStudy (anonymous):

Dexter bought a glass pyramid from a gift store. The pyramid, with a height of 5 cm and a square base of side 3 cm, was packed in a box shaped as a cube of side 6 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue?

OpenStudy (istim):

Isn't it just the volume of the cube subtracted by the volume of the pyramid then?

OpenStudy (istim):

In such case the formula for the cube's volume is v=lwh, but I don't know what the formula would be for the square-based pyramid.

OpenStudy (anonymous):

I am not really sure.. but i think the answer is 201... ? Is that correct?

OpenStudy (istim):

I don't know. Haven't done the math myself.

OpenStudy (anonymous):

can you do it? lol

OpenStudy (istim):

I'm a bit busy myself. Have confidence. Try re-doing it without looking at your original work if you truly are concerned.

OpenStudy (istim):

Oh, volume for square-based pyramid is v=1/3b^2h

OpenStudy (istim):

I'm trying to help.

OpenStudy (anonymous):

I know. thanks.. ill try with the formula u gave me..

OpenStudy (istim):

So I guess your complete formula for the paper tissue is v=lwh -1/3b^2h

OpenStudy (anonymous):

ahh i don't have the length though.

OpenStudy (istim):

Which one?

OpenStudy (istim):

The cube? All lengths are the same. It's like s^3.

OpenStudy (anonymous):

of the pyramid.... i need lwh don't i?

OpenStudy (istim):

LWh is the cube...

OpenStudy (anonymous):

ohh. so its all the same??

OpenStudy (istim):

What's the same?

OpenStudy (istim):

Are you good then?

OpenStudy (anonymous):

im so confused to be honest..

OpenStudy (istim):

Hello?

OpenStudy (istim):

What's the confusion? Just read their our chat to understand...

OpenStudy (anonymous):

what??

OpenStudy (istim):

I don't see what the confusion is.

OpenStudy (anonymous):

how to get the answer. if you don't understand what the confusion is.. can you help me, and give me an answer! pleassseee.

OpenStudy (istim):

I gave you the combined formula, v=lwh -1/3b^2h. Just plug and play.

OpenStudy (anonymous):

i did. but i don't have all the things needed for the equation!!!

OpenStudy (istim):

The values themselves are given in your question.

OpenStudy (anonymous):

not all of them..

OpenStudy (istim):

lwh is the cube. The other one is the square-based pyramid.

OpenStudy (anonymous):

can you help me plug them in ? please

OpenStudy (istim):

l,w and h are all 6.

OpenStudy (anonymous):

okay. so v=lwh -1/3b^2h, v=(6)(6)(6)- 1/3(5^2)(3)

OpenStudy (istim):

Because it is a cube. Cubes have side lengths which are all the same.

OpenStudy (anonymous):

are you sure the pyramid formula is squared?

OpenStudy (istim):

It's not squared. Ignore that if you don't understand.

OpenStudy (anonymous):

ohh. okay. so then its just v=(6)(6)(6)- 1/3(5)(3), 216-5=209?

OpenStudy (anonymous):

i mean.. 211?

OpenStudy (istim):

I don't know actually. I haven't done the calculations myself.

OpenStudy (anonymous):

Woah. I'm sorry. thanks you really helped! I got the question CORRECT:) anddd can you help me with my other questions.

OpenStudy (anonymous):

Ross is constructing two similar boxes shaped as similar rectangular prisms. The larger box has a height of 20 inches and a square base of side length 12 inches. If the smaller box has a height of 15 inches, what is its volume?

OpenStudy (istim):

Uh. Maybe ask the other guys.

OpenStudy (anonymous):

lol. kk